# Sequential Monte Carlo with Gaussian Mixture Approximation for Infinite-Dimensional Statistical Inverse Problems

### v3

@@ 2.2 Approximation theory of SMC
$\begin{split}\frac{\mu^{\mathrm{d}}(dv)Q(v,du)}{\mu^{\mathrm{d}}(du)Q(u,dv)}=&\frac{\mu^{\mathrm{d}}(dv)Q(v,du)}{\mu_{0}(dv)Q(v,du)}\frac{\mu_{0}(dv)Q(v,du)}{\mu_{0}(du)Q(u,dv)}\frac{\mu_{0}(du)Q(u,dv)}{\mu^{\mathrm{d}}(du)Q(u,dv)},\\ =&e^{\Phi(u)-\Phi(v)}\frac{\mu_{0}(dv)Q(v,du)}{\mu_{0}(du)Q(u,dv)}\end{split}$ (2.34)
@@
1:Input: Specify the prior $\mu_{0}$. Draw $u_{0}\sim\mu_{0}.$.
@@ 2.4 Approximation theory of Gaussian mixtures
In SMC, we need to choose a transition kernel in the mutation step, and the simplest choice is to sample from the prior, i.e., to use an independence sampler. This sampler method works well when the likelihood is not too informative, but works poorly if the information in the likelihood is substantial Dashti2017 . To utilize the information from the likelihood, several proposals are available, for example, the pCN, the pCN Langevin Andrew_review , and the Hessian-preconditioned explicit Langevin CUI2016 , each utilizing information of $\Phi$, $D\Phi$, and $D^{2}\Phi$ respectively. We need to solve one, two, and four PDEs, respectively, to compute them with the adjoint method Ghattas2021LearningPM , which is computationally intensive. To utilize likelihood information while avoiding the computation of the likelihood function, we use a Gaussian mixture approximation of the posterior. In other words, we retain the first step of pCN-GM, which proposes a sample, and discard the second step of the acceptance process, accepting all proposals directly. The resulting SMC method is SMC-GM, because the mutation step is simply a Gaussian mixture sampler. SMC-GM makes sense only if we can find a good Gaussian mixture approximation to the posterior. In this subsection, we prove the denseness of the Gaussian mixture measure in the posterior measures, provide an error estimate, and propose the SMC-GM method finally.
@@
Then it follows from Theorem 5.(d) in Nguyen2020approximation that Gaussian mixtures have the universal approximation property in the finite-dimensional setting. If $f$ and $g$ are probability density functions and $g\in\mathcal{C}_{0}$, then for any measurable $f$, there exists a sequence $\{h_{M}\}\in\mathcal{M}_{M}^{g}$ such that
@@
The above theorem states that for posterior measures in a separable Hilbert space, Gaussian mixture measures are dense within them; furthermore, this Gaussian mixture is also mutually equivalent to the prior measure. This theorem is the motivation for our use of Gaussian mixtures for approximation. The proof is given in section 5.4.7 in the supplementary material.
In practice, we estimate the parameters of a Gaussian mixture from samples using the method from fengzhe2018 , which is established in function space, and the Gaussian mixture is also equivalent to the prior measure $\mu_{0}$. Then we show that sampling from Gaussian mixture is an approximation of pCN-GM. Let $\beta=1$ in (2.23), and we obtain
@@
The Gaussian mixture sampler $Q$ is an approximation of the transition kernel $P$, defined by (2.17):
$d(Q\mu^{N},P\mu^{N})<2d(Q,\mu_{post})+2d(\mu^{N},\mu_{post}).$ (2.41)
The proof is given in section 5.4.8 in the supplementary material. This theorem indicates that if we have a good Dirac approximation of the posterior $\mu^{N}$, and a good Gaussian mixture approximation of the posterior $Q$, then we can directly use Gaussian mixture sampler to replace the transition kernel $P$. The resulting SMC-GM is concluded in Algorithm 3.
@@ 3 Numerical Examples
In this section, we provide three numerical examples to substantiate the results in Section 2. We apply the SMC-GM method to two typical inverse problems, including one linear and one nonlinear problem. For the nonlinear problem, we also consider a sparse measurement situation as the third example. Through these experiments, we demonstrate that the SMC-GM algorithm is faster than SMC-pCN and has strong sampling capabilities for multi-modal posteriors. In the following, SMC-pCN is used for comparison. Thus we briefly discuss the computational cost of them. If we require $N$ samples, and the number of intermediate measures is $J$, then SMC-GM method only need to compute the weights in the resampling step for each particle, and a total of $NJ$ PDEs. As for SMC-pCN method, we need to sample a Markov chain for every particle in each layer, the length of the chain can be chosen between 5 and 1000 Beskos2015SC . In all three numerical examples, we will select Markov chains of length 200, requiring us to solve a total of $200NJ$ PDEs. Note that the method for determining $h_{i}$ is described in section 5.3 in the supplementart material and we estimate the parameters of a Gaussian mixture from samples using the method from fengzhe2018 .
@@ 3.1 A multi-modal example
As the first example, we construct a multi-modal problem in a 1-dimensional interval $\Omega=[0,1]$ as described in fengzhe2018 . Let the unknown $u\in\mathcal{H}=L^{2}(\Omega)$ and assume the prior is a zero-mean Gaussian measure on $\mathcal{H}$ with covariance operator $\mathcal{C}=(I-0.01\Delta)^{-2}$. We consider a four-modal likelihood function given by
@@ 3.2 Elliptic PDE Inverse Problem
    The initial step norm is $\beta=0.2$, then we update it using an adaptive strategy, as stated in section 5.3 in the supplementary material.
@@
where $\overline{u}(x)$ is the mean function, $e_{k}$ is the $k$-th basis function, and $u_{k}$ is the $k$-th Fourier coefficient. Due to Theorem VI.16 in reed1981functional , the eigenfunctions of the prior covariance operator form a complete orthonormal basis. Thus, we can choose them as $e_{k}$.
@@
Finally, we show the mesh independence of the SMC-GM method, a key property for a well-defined function space method. Figure 4(a) shows temperature plots for discrete dimensions $N=\{400,1600,3600,6400,10000\}$, starting at 0 and increasing to 1. We observe that the temperatures show mesh independence.
@@ 4 Discussion
In this paper, we propose a novel transition kernel, pCN-GM, which is derived from the Crank-Nicolson discretization of the Langevin system and is proven to be well-defined in infinite-dimensional spaces. As a theoretical foundation for its application in SMC, we prove an SMC convergence theorem under weaker conditions. Given the proven density of the Gaussian mixture measure, we propose an efficient approximation algorithm called SMC-GM and apply it to three typical inverse problems. All numerical results indicate that SMC-GM excels over SMC-pCN in terms of efficiency and multi-modal sampling capability, with comparable accuracy, and also exhibits mesh-independence.
We have not provided an explicit expression for the error introduced by the Gaussian mixture approximation in the infinite-dimensional setting, which has been studied in the finite-dimensional context FengshuoZhang2020 . A possible further algorithmic innovation is to use multi-level Monte Carlo method MLMC within SMC.
@@ 5.2 From Langevin system to pCN-GM proposal
Discretize the equation using the Crank-Nicolson method, it follows that
@@ 5.3 Some numerical details
Determine the temperatures: We need to choose the temperatures $0=h_{1}\leq h_{2}\leq\cdots\leq h_{J-1}<h_{J}=1$. If these temperatures are too dense, the SMC will run slowly; if they are too sparse, the adjacent posteriors $\mu_{j}$ and $\mu_{j+1}$ will differ significantly, which is not conducive to importance sampling Dai2022 . We can use the effective sample size (ESS) to select appropriate temperatures:
@@
where $w_{i}$ is the weight of the $i$-th particle. A small $h_{j}$ can result in a large ESS. An excessively large value of $h$ can cause the loss of a large number of samples during the resampling step, thereby leading to a small ESS. Thus, we can use a simple bisection method to find an $h_{j}$ such that $\text{ESS}>N_{\text{thresh}}\mathrel{\mathop{\ordinarycolon}}=0.6N$ Beskos2015SC .
@@ 5.4.1 Discussion of total variation distance
Under Assumptions 2.1, let $\nu_{j}=P_{j}\mu_{j},$ then we have
$d(L_{j}\nu_{j-1},L_{j}\mu)\leq 2\kappa_{1}^{-1}\kappa_{2}^{-1}d(\nu_{j-1},\mu).$
Let $g_{j}(v)\mathrel{\mathop{\ordinarycolon}}=e^{-\Phi_{j}(u)}$, then it follows from Assumptions 2.1(a) that $|\kappa_{1}g_{j}(v)|\leq 1$. For simplicity we denote $g^{\prime}=\kappa_{1}g_{j}$ with $|g^{\prime}|\leq 1$. Moreover, with Assumptions 2.1(c) we know that
$\begin{split}\nu_{j-1}(g_{j})=&\int g_{j}P_{j-1}\mu_{j-1}(dm)\\ =&\int g_{j}\mu_{j-1}(dm)\\ =&\int e^{-\Phi_{j}(m)}e^{-\Phi_{1}(m)-\dots-\Phi_{j-1}(m)}\mu_{0}(dm)\\ \geq&\int e^{-\sum_{j=1}^{J}\Phi_{j}(m)}\mu_{0}(dm)\\ =&\int e^{-\Phi(m)}\mu_{0}(dm)\\ >&\kappa_{2}.\end{split}$ (5.4)
The aim is to measure the distance between $L_{j}\nu_{j-1}$ and $L_{j}\mu$. Hence we need to compute the difference between the integral of $f$ with respect to two measures, i.e.,
$(L_{j}\nu_{j-1})(f)-(L_{j}\mu)(f)$ $=\frac{\nu_{j-1}(fg_{j})}{\nu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\mu(g_{j})}$
$=\frac{\nu_{j-1}(fg_{j})}{\nu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\nu_{j-1}(g_{j})}+\frac{\mu(fg_{j})}{\nu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\mu(g_{j})}$
$=\frac{\kappa_{1}^{-1}}{\nu_{j-1}(g_{j})}[\nu_{j-1}(fg^{\prime})-\mu(fg^{\prime})]+\frac{\mu(fg_{j})}{\mu(g_{j})}\frac{\kappa_{1}^{-1}}{\nu_{j-1}(g_{j})}[\mu(g^{\prime})-\nu_{j-1}(g^{\prime})].$
@@
$|(L_{j}\nu_{j-1})(f)-(L_{j}\mu)(f)|^{2}\leq$ $\frac{2\kappa_{1}^{-2}}{\nu_{j-1}(g_{j})^{2}}|\nu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}$
$+2\frac{\mu(fg_{j})^{2}}{\mu(g_{j})^{2}}\frac{\kappa_{1}^{-2}}{\nu_{j-1}(g_{j})^{2}}|\mu(g^{\prime})-\nu_{j-1}(g^{\prime})|^{2}.$
By applying inequality (5.4), we have
$|(L_{j}\nu_{j-1})(f)-(L_{j}\mu)(f)|^{2}\leq 2\kappa_{1}^{-2}\kappa_{2}^{-2}|\nu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2|f|_{\infty}^{2}\kappa_{1}^{-2}\kappa_{2}^{-2}|\mu(g^{\prime})-\nu_{j-1}(g^{\prime})|^{2}.$
@@
$d(L_{j}\nu_{j-1},L_{j}\mu)$
$=\sup_{|f|_{\infty}\leq 1}\sqrt{\mathbb{E}_{\omega}|(L_{j}\nu_{j-1})(f)-(L_{j}\mu)(f)|^{2}}$
$\leq\sup_{|f|_{\infty}\leq 1}\sqrt{2\kappa_{1}^{-2}\kappa_{2}^{-2}\mathbb{E}_{\omega}|\nu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2|f|_{\infty}^{2}\kappa_{1}^{-2}\kappa_{2}^{-2}\mathbb{E}_{\omega}|\mu(g^{\prime})-\nu_{j-1}(g^{\prime})|^{2}}$
$\leq\sqrt{2\kappa_{1}^{-2}\kappa_{2}^{-2}\sup_{|f|_{\infty}\leq 1}\mathbb{E}_{\omega}|\nu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2\kappa_{1}^{-2}\kappa_{2}^{-2}d(\mu,\nu_{j-1})^{2}}$
$=\sqrt{2\kappa_{1}^{-2}\kappa_{2}^{-2}d(\mu,\nu_{j-1})^{2}+2\kappa_{1}^{-2}\kappa_{2}^{-2}d(\mu,\nu_{j-1})^{2}}$
$=\frac{2}{\kappa_{1}\kappa_{2}}d(\mu,\nu_{j-1}),$
@@
5.4.3 Proof of Theorem 2.2 (convergence of SMC)
@@
$(\text{triangle inequality})\quad$ $\leq 2\kappa_{1}^{-1}\kappa_{2}^{-1}[d(S^{N}P_{j}\mu_{j}^{N},P_{j}\mu_{j}^{N})+d(P_{j}\mu_{j}^{N},\mu_{j})]$

### v4

@@ 2.2 Approximation theory of SMC
$\begin{split}\frac{\rho(dv)Q(v,du)}{\rho(du)Q(u,dv)}=&\frac{\rho(dv)Q(v,du)}{\mu_{0}(dv)Q(v,du)}\frac{\mu_{0}(dv)Q(v,du)}{\mu_{0}(du)Q(u,dv)}\frac{\mu_{0}(du)Q(u,dv)}{\rho(du)Q(u,dv)},\\ =&\frac{d\rho/d\mu_{0}(v)}{d\rho/d\mu_{0}(u)}\frac{\mu_{0}(dv)Q(v,du)}{\mu_{0}(du)Q(u,dv)}\end{split}$ (2.34)
@@
1:Input: Specify the prior $\mu_{0}$. Draw $u_{0}\sim\mu_{0}.$
@@
Using Algorithm 2 as the mutation step in Algorithm 1 yields the SMC-pCN-GM algorithm. The approximation properties of SMC-pCN-GM are established in Theorem 2.2, while its well-definedness is ensured by Theorem 2.4.
@@ 2.4 Approximation theory of Gaussian mixtures
In SMC, we need to choose a transition kernel in the mutation step, and the simplest choice is to sample from the prior, i.e., to use an independence sampler. This sampler method works well when the likelihood is not too informative, but works poorly if the information in the likelihood is substantial Dashti2017 . To utilize the information from the likelihood, several proposals are available, for example, the pCN, the pCN Langevin Andrew_review , and the Hessian-preconditioned explicit Langevin Cui2016 , each utilizing information of $\Phi$, $D\Phi$, and $D^{2}\Phi$ respectively. We need to solve one, two, and four PDEs, respectively, to compute them with the adjoint method Ghattas2021 , which is computationally intensive. To utilize likelihood information while avoiding the computation of the likelihood function, we use a Gaussian mixture approximation of the posterior. In other words, we retain the first step of pCN-GM, which proposes a sample, and discard the second step of the acceptance process, accepting all proposals directly. The resulting SMC method is SMC-GM, because the mutation step is simply a Gaussian mixture sampler. SMC-GM makes sense only if we can find a good Gaussian mixture approximation to the posterior. In this subsection, we prove the denseness of the Gaussian mixture measure in the posterior measures, provide an error estimate, and propose the SMC-GM method finally.
@@
Then it follows from Theorem 5.(d) in Nguyen2020 that Gaussian mixtures have the universal approximation property in the finite-dimensional setting. If $f$ and $g$ are probability density functions and $g\in\mathcal{C}_{0}$, then for any measurable $f$, there exists a sequence $\{h_{M}\}\in\mathcal{M}_{M}^{g}$ such that
@@
The above theorem states that for posterior measures in a separable Hilbert space, Gaussian mixture measures are dense within them; furthermore, this Gaussian mixture is also mutually equivalent to the prior measure. This theorem is the motivation for our use of Gaussian mixtures for approximation. The proof is given in 5.4.8.
In practice, we estimate the parameters of a Gaussian mixture from samples using the method from Feng2018 , which is established in function space, and the Gaussian mixture is also equivalent to the prior measure $\mu_{0}$. Then we show that sampling from Gaussian mixture is an approximation of pCN-GM. Let $\beta=1$ in (2.23), and we obtain
@@
Recall that $\mu_{j}$ is the posterior of layer j. Then we have that the Gaussian mixture sampler $Q$ is an approximation of the transition kernel $P$, defined by (2.17):
$d(Q\mu_{j}^{N},P\mu_{j}^{N})\leq d(Q,\mu_{j})+d(\mu_{j}^{N},\mu_{j}).$ (2.41)
Moreover there exists a Gaussian mixture measure $Q$ such that
$d^{\prime}(Q\mu_{j}^{N},P\mu_{j}^{N})\leq(2+\epsilon)d^{\prime}(\mu_{j}^{N},\mu_{j}),\text{ for any }\epsilon>0,$ (2.42)
where $d^{\prime}$ is a weaker distance defined by
$d^{\prime}(\mu,\nu)\mathrel{\mathop{\ordinarycolon}}=\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\sqrt{\mathbb{E}\Big[\int fd\mu-\int fd\nu\Big]^{2}}.$
Note that the original error bound still holds, since $d^{\prime}(\mu,\nu)\leq d(\mu,\nu)$. The proof of Theorem 2.6 is given in 5.4.9. This theorem indicates that, if we have a good Dirac approximation of the posterior, then it is reasonable to directly use a Gaussian mixture sampler to replace the transition kernel $P$. The resulting SMC-GM is concluded in Algorithm 3.
@@ 3 Numerical Examples
In this section, we provide three numerical examples to substantiate the results in Section 2. Compared with SMC-pCN-GM, the reduced SMC-GM algorithm requires fewer computational resources while maintaining satisfactory performance. Hence, we mainly report the numerical results of the SMC-GM algorithm. We apply the SMC-GM method to two typical inverse problems, including one linear and one nonlinear problem. For the nonlinear problem, we also consider a sparse measurement situation as the third example. Through these experiments, we demonstrate that the SMC-GM algorithm is faster than SMC-pCN and has strong sampling capabilities for multi-modal posteriors. In the following, SMC-pCN is used for comparison. Thus we briefly discuss the computational cost of them. If we require $N$ samples, and the number of intermediate measures is $J$, then SMC-GM method only need to compute the weights in the resampling step for each particle, and a total of $NJ$ PDEs. As for SMC-pCN method, we need to sample a Markov chain for every particle in each layer, the length of the chain can be chosen between 5 and 1000 Beskos2015SC . In all three numerical examples, we will select Markov chains of length 200, requiring us to solve a total of $200NJ$ PDEs. Note that the method for determining $h_{i}$ is described in section 5.3 in the supplementart material and we estimate the parameters of a Gaussian mixture from samples using the method from Feng2018 . All the programming codes are accessible at https://github.com/jjx323/IPBayesML-SMC-GM. SMC is intrinsically parallel, and so are our implementations.
@@ 3.1 A multi-modal example
As the first example, we construct a multi-modal problem in a 1-dimensional interval $\Omega=[0,1]$ as described in Feng2018 . Let the unknown $u\in\mathcal{H}=L^{2}(\Omega)$ and assume the prior is a zero-mean Gaussian measure on $\mathcal{H}$ with covariance operator $\mathcal{C}=(I-0.01\Delta)^{-2}$. We consider a four-modal likelihood function given by
@@ 3.2 Elliptic PDE Inverse Problem
    The initial step norm is $\beta=0.2$, then we update it using an adaptive strategy (see 5.3).
@@
where $\overline{u}(x)$ is the mean function, $e_{k}$ is the $k$-th basis function, and $u_{k}$ is the $k$-th Fourier coefficient. Due to Theorem VI.16 in Reed1980 , the eigenfunctions of the prior covariance operator form a complete orthonormal basis. Thus, we can choose them as $e_{k}$.
@@
Finally, we show the mesh independence of the SMC-GM method, a key property for a well-defined function space method. At each layer, the particle system determines the temperature of the next layer through the procedure described in 5.3. As a result, differences between particle systems arising from different discretizations are reflected through the corresponding temperatures. Figure 4(a) shows temperature plots for discrete dimensions $N=\{400,1600,3600,6400,10000\}$, starting at 0 and increasing to 1. We observe that the temperatures show mesh independence.
@@ 4 Discussion
In this paper, we propose a novel transition kernel, pCN-GM, which is derived from the Crank-Nicolson discretization of the Langevin system and is proven to be well-defined in infinite-dimensional spaces. By employing pCN-GM in the mutation step of SMC, we obtain the SMC-pCN-GM algorithm. As a theoretical foundation for SMC, we prove a convergence theorem under weaker conditions. Given the proven density of the Gaussian mixture measure, we propose an efficient approximation algorithm called SMC-GM and apply it to three typical inverse problems. All numerical results indicate that SMC-GM excels over SMC-pCN in terms of efficiency and multi-modal sampling capability, with comparable accuracy, and also exhibits mesh-independence.
We have not provided an explicit expression for the error introduced by the Gaussian mixture approximation in the infinite-dimensional setting, which has been studied in the finite-dimensional context Zhang2020 . Further acceleration of SMC is an interesting direction that we plan to explore in future work. One possible approach is to introduce parallel MCMC schemes into the mutation step of SMC, for example via multi-level Monte Carlo methods Wang2023 . Another possibility is to parallelize SMC itself Verge2015 , by running two or more SMC samplers simultaneously and allowing state exchange between them.
@@ 5.1 Four transition kernels
We then discuss whether the pCN algorithm is irreducible in the infinite-dimensional Hilbert space setting, and its implications for the SMC framework. In the infinite-dimensional setting, the proposal of pCN is
$P(u,dv)=\sqrt{1-\beta^{2}}u+\beta^{2}\xi,\quad\xi\sim\mu_{prior}=\mathcal{N}(0,\mathcal{C}),u,v\in\mathcal{H},$
and $\mathcal{C}$ is a positive, symmetric, trace-class operator. In infinite-dimensional spaces, the Lebesgue measure does not exist, and two Gaussian measures with different means are mutually singular if the difference of their means lies outside the Cameron–Martin space $\tilde{\mathcal{H}}$. The singularity of $\mu$ and $\nu$ means that there exists a measurable set $A$ such that $\mu(A)=1$ and $\nu(A)=0$, which is denoted by $\mu\perp\nu$. Thus we have
$\mathcal{N}(0,\mathcal{C})\perp\mathcal{N}(\sqrt{1-\beta^{2}}x,\beta^{2}\mathcal{C}),\quad\forall x\notin\tilde{\mathcal{H}}$.
As a consequence of this singularity, for sets $A$ with positive posterior measure $\mu_{\text{post}}(A)>0$, it generally holds that $P(u,A)=0$ for $\mu_{\text{post}}$-almost all $u$ Hairer2014 . Therefore, under the standard finite-dimensional definition of irreducibility, the pCN transition kernel on infinite-dimensional state spaces fails to be irreducible. This phenomenon is not due to the properties of the pCN algorithm, but rather to the intrinsic nature of the infinite-dimensional setting.
Consequently, alternative approaches are required to establish ergodicity and related results. It is well known that for Markov chains with reversible transition kernels, a strictly positive spectral gap is equivalent to a form of geometric ergodicity Roberts1997 . The uniqueness of the invariant distribution of pCN, together with the existence of an $L^{2}$-spectral gap and a Wasserstein spectral gap, is established in Hairer2014 . Moreover, this spectral gap is dimension-independent, and laws of large numbers as well as central limit theorems are derived. Therefore, using pCN to sample in infinite-dimensional function spaces is feasible.
Finally, in SMC methods, the primary sampling mechanism is importance sampling, while MCMC steps are introduced to enhance particle diversity. Neither the theoretical analysis nor practical implementations of SMC require the Markov transitions to be run for long time horizons, and irreducibility of the underlying Markov kernels is therefore not a central requirement within the SMC framework.
@@ 5.2 From Langevin system to pCN-GM proposal
Discretize the equation using the Crank-Nicolson method, and it follows that
@@ 5.3 Some numerical details
Determine the temperatures: We need to choose the temperatures $h_{j}\in(0,1),1\leq i\leq J$ such that $\sum_{j=1}^{J}h_{j}=1$. If these temperatures are too small, the SMC will run slowly; if they are too large, the adjacent posteriors $\mu_{j}$ and $\mu_{j+1}$ will differ significantly, which is not conducive to importance sampling Dai2022 . We can use the effective sample size (ESS) to select appropriate temperatures:
@@
where $w_{i}$ is the weight of the $i$-th particle. A small $h_{j}$ can result in a large ESS. An excessively large value of $h_{j}$ can cause the loss of a large number of samples during the resampling step, thereby leading to a small ESS. Thus, we can use a simple bisection method to find an $h_{j}$ such that $\text{ESS}>N_{\text{thresh}}\mathrel{\mathop{\ordinarycolon}}=0.6N$ Beskos2015SC .
@@ 5.4.1 Discussion of total variation distance
Under Assumptions 2.1, we have
$d(L_{j}\mu_{j-1},L_{j}\mu)\leq 2\kappa_{1}^{-1}\kappa_{2,j-1}^{-1}d(\mu_{j-1},\mu).$
Let $g_{j}(v)\mathrel{\mathop{\ordinarycolon}}=e^{-\Phi_{j}(u)}$, then it follows from Assumptions 2.1(a) that $|\kappa_{1}g_{j}(v)|\leq 1$. For simplicity we denote $g^{\prime}=\kappa_{1}g_{j}$ with $|g^{\prime}|\leq 1$. The aim is to measure the distance between $L_{j}\mu_{j-1}$ and $L_{j}\mu$. Hence we need to compute the difference between the integral of $f$ with respect to two measures, i.e.,
$(L_{j}\mu_{j-1})(f)-(L_{j}\mu)(f)$ $=\frac{\mu_{j-1}(fg_{j})}{\mu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\mu(g_{j})}$
$=\frac{\mu_{j-1}(fg_{j})}{\mu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\mu_{j-1}(g_{j})}+\frac{\mu(fg_{j})}{\mu_{j-1}(g_{j})}-\frac{\mu(fg_{j})}{\mu(g_{j})}$
$=\frac{\kappa_{1}^{-1}}{\mu_{j-1}(g_{j})}[\mu_{j-1}(fg^{\prime})-\mu(fg^{\prime})]+\frac{\mu(fg_{j})}{\mu(g_{j})}\frac{\kappa_{1}^{-1}}{\mu_{j-1}(g_{j})}[\mu(g^{\prime})-\mu_{j-1}(g^{\prime})].$
@@
$|(L_{j}\mu_{j-1})(f)-(L_{j}\mu)(f)|^{2}\leq$ $\frac{2\kappa_{1}^{-2}}{\mu_{j-1}(g_{j})^{2}}|\mu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}$
$+2\frac{\mu(fg_{j})^{2}}{\mu(g_{j})^{2}}\frac{\kappa_{1}^{-2}}{\mu_{j-1}(g_{j})^{2}}|\mu(g^{\prime})-\mu_{j-1}(g^{\prime})|^{2}.$
By applying Assumptions 2.1, we have
$|(L_{j}\mu_{j-1})(f)-(L_{j}\mu)(f)|^{2}\leq 2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}|\mu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2|f|_{\infty}^{2}\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}|\mu(g^{\prime})-\mu_{j-1}(g^{\prime})|^{2}.$
@@
$d(L_{j}\mu_{j-1},L_{j}\mu)$
$=\sup_{|f|_{\infty}\leq 1}\sqrt{\mathbb{E}_{\omega}|(L_{j}\mu_{j-1})(f)-(L_{j}\mu)(f)|^{2}}$
$\leq\sup_{|f|_{\infty}\leq 1}\sqrt{2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}\mathbb{E}_{\omega}|\mu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2|f|_{\infty}^{2}\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}\mathbb{E}_{\omega}|\mu(g^{\prime})-\mu_{j-1}(g^{\prime})|^{2}}$
$\leq\sqrt{2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}\sup_{|f|_{\infty}\leq 1}\mathbb{E}_{\omega}|\mu_{j-1}(fg^{\prime})-\mu(fg^{\prime})|^{2}+2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}d(\mu,\mu_{j-1})^{2}}$
$=\sqrt{2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}d(\mu,\mu_{j-1})^{2}+2\kappa_{1}^{-2}\kappa_{2,j-1}^{-2}d(\mu,\mu_{j-1})^{2}}$
$=\frac{2}{\kappa_{1}\kappa_{2,j-1}}d(\mu,\mu_{j-1}),$
@@
5.4.3 Proof of Lemma 2.1(d) (error of operator $R$)
Let $\{v^{n}\}_{n=1}^{N}$ be the samples from $\mu$. After reweighting, we have a new measure
$\tilde{\mu}^{N}=\sum_{n=1}^{N}w_{n}(v^{n})\,\delta_{v^{n}}.$
Resampling produces new samples $\tilde{v}^{n}\sim\tilde{\mu}^{N}$ (conditionally independent given $\{v^{n}\}_{n=1}^{N}$), which define the empirical measure at layer $j$:
$\mu^{N}=\frac{1}{N}\sum_{i=1}^{N}\delta(v-\tilde{v}^{i}).$
The following identity is useful:
$\mathbb{E}_{v\sim\mu^{N}}[f(v)\mid\{v^{n}\}_{n=1}^{N}]=$ $\int f(v)\frac{1}{N}\sum_{n=1}^{N}\delta(v-v^{i})\,dv$
$=$ $\frac{1}{N}\sum_{i=1}^{N}f(\tilde{v}^{i})=\mu^{N}(f).$
We now compute the resampling error:
$\mathbb{E}_{\tilde{v}^{n}}\bigl|\tilde{\mu}^{N}(f)-\mu^{N}(f)\bigr|^{2}=\mathbb{E}_{v^{n}}\Bigl[\mathbb{E}_{\tilde{v}^{n}\sim\tilde{\mu}^{N}\mid v^{n}}\bigl[(\tilde{\mu}^{N}(f)-\mu^{N}(f))^{2}\mid\{v^{i}\}_{i=1}^{N}\bigr]\Bigr]=\mathbb{E}\!\left[\operatorname{Var}\bigl(\tilde{\mu}^{N}(f)\mid\{v^{i}\}_{i=1}^{N}\bigr)\right].$
The inner expectation is with respect to the randomness of the resampling procedure itself, while the outer expectation is taken over the samples $\{v^{n}\}$. The final equality follows since
$\mathbb{E}_{\tilde{v}}\bigl[\tilde{\mu}^{N}(f)\mid\{v^{i}\}_{i=1}^{N}\bigr]=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\tilde{v}}[f(\tilde{v}^{i})\mid\{v^{i}\}_{i=1}^{N}]=\mathbb{E}[f(\tilde{v}^{i})\mid\{v^{i}\}_{i=1}^{N}]=\mu^{N}(f).$
We compute the conditional variance:
$\operatorname{Var}\bigl(\tilde{\mu}^{N}(f)\mid\{v^{i}\}_{i=1}^{N}\bigr)$ $=\operatorname{Var}\!\left(\frac{1}{N}\sum_{n=1}^{N}f(\tilde{v}^{n})\Bigm|\mu^{N}\right)$
$=\frac{1}{N^{2}}\sum_{n=1}^{N}\operatorname{Var}\bigl(f(\tilde{v}^{n})\mid\{v^{i}\}_{i=1}^{N}\bigr)$
$=\frac{1}{N}\mathbb{E}\bigl[(f(\tilde{v}^{n})-\mu^{N}(f))^{2}\mid\{v^{i}\}_{i=1}^{N}\bigr]$
$=\frac{1}{N}\Bigl(\mu^{N}(f^{2})-\mu^{N}(f)^{2}\Bigr).$
Consequently,
$\mathbb{E}\bigl|\tilde{\mu}^{N}(f)-\mu^{N}(f)\bigr|^{2}$ $=\frac{1}{N}\,\mathbb{E}\!\left[\mu^{N}(f^{2})-\mu^{N}(f)^{2}\right]$
$=\frac{1}{N}\,\mathbb{E}\left[\sum_{i=1}^{N}w_{i}f(v^{i})^{2}-\Bigl(\sum_{i=1}^{N}w_{i}f(v^{i})\Bigr)^{2}\right].$
Here $\mathbb{E}$ is taken with respect to the samples $\{v^{i}\}$. Since both the weights $w_{i}$ and the values $f(v^{i})$ depend on $v^{i}$, this expectation is difficult to evaluate explicitly. However, taking the supremum over all test functions satisfying $\|f\|_{\infty}\leq 1$, we immediately obtain
$\mathbb{E}\bigl|\tilde{\mu}^{N}(f)-\mu^{N}(f)\bigr|^{2}$ $\leq\frac{1}{N},$
$d(\tilde{\mu}^{N},\mu^{N})$ $\leq\frac{1}{\sqrt{N}},$
which represents the resampling error. Let $R$ denote the resampling operator, then it follows that
$d(R\mu^{N},\mu^{N})\leq\frac{1}{\sqrt{N}}.$
5.4.4 Proof and discussion of Theorem 2.2 (convergence of SMC)